OpenStudy (anonymous):

How would this problem be solved? D=1/4(e-3f) for f

OpenStudy (anonymous):

multiply by 4, subtract 3 divide by -3 ta da

OpenStudy (anonymous):

how would it be worked out?

OpenStudy (anonymous):

sorry subtract e

OpenStudy (anonymous):

can you try what i said first , before I give an answer

OpenStudy (anonymous):

D=1/4(e-3f) for f Distribute 4 D = 1/(4e - 12f) Multiply the denominator so you can cancel it. D(4e - 12f) = 1 Distribute D 4eD - 12fD = 1 Isolate f -12fD = 1 - 4eD 12fD = 4eD - 1 f = (4eD - 1)/12D

OpenStudy (anonymous):

wow over complicated much?

OpenStudy (anonymous):

so compilcated but thanks yallll

OpenStudy (anonymous):

multiply by 4 4D=e-3f subtract e 4d-e=-3f divide by -3 (4d-e)/-3=f

OpenStudy (anonymous):

do you want me to show you with more instruction?

OpenStudy (anonymous):

opps I was wrong, it happens, sorry @franciscanmonk WAS CORRECT